# Linking Potentially Misclassified Healthy Food Access to Diabetes Prevalence

### v1

@@ 1 Introduction
Maintaining a healthy diet is more than simply a choice. It requires a person to have consistent access to and ability to obtain healthy food. Physical access to healthy foods may be hindered by many factors, including the neighborhood’s geography or community’s socioeconomic makeup (Wolfson et al., 2019). Subsequently, neighborhoods with poor food access may see higher prevalences of diseases linked to healthy eating (Gucciardi et al., 2014). We focus on the relationship between diabetes prevalence and food access at the neighborhood (census tract) level, although our methods adapt to other outcomes as well.
@@
To estimate this relationship, we must first quantify food access, for which there are multiple metrics. We devote our attention to “access indicators,” which classify neighborhoods into binary categories of having/not having access to healthy food within some fixed distance. The United States Department of Agriculture (USDA) includes indicators like these in their Food Research Atlas (U.S. Department of Agriculture (USDA), 2019). Computing these food access indicators requires researchers to first calculate the pairwise distances between neighborhoods and potential healthy food retailers. Just like there are multiple ways to measure food access, there are multiple ways to measure these distances. For instance, the Haversine formula modifies the traditional straight-line Euclidean distance between two points (e.g., a neighborhood centroid and grocery store) to account for the curvature of the Earth. Alternatively, route-based distance calculations are possible, which only allow travel via established infrastructure (like roads) to connect the two points.
@@
These methods each have their advantages, and choosing between them to calculate distances requires a trade-off. The Haversine formula is a simple mathematical equation, making it free and easy to calculate in existing software (e.g., Hijmans, 2022). However, it underestimates distances from neighborhoods to healthy food retailers and therefore overestimates neighborhoods’ food access. This underestimation stems from the fact that straight-line distances are necessarily less than or equal to any other measure of distance. In our study, underestimated Haversine distances can lead to neighborhoods being misclassified as nominally having access to healthy foods when they truly do not (false positives). Specifically, the access indicators are subject to one-sided misclassification, since false positives are the only possible type of error.
@@
In short, Haversine distance is easily obtained but leads to access indicators that may be incorrect. Taking route-based distances via software like Google Maps circumvents the possibility of unrealistic, impassable routes. Still, these more-accurate distances are computationally expensive and, despite availability of open source options, may also be financially expensive. Therefore, calculating them for all neighborhoods may not be plausible.
@@
For our study, with the prevalence of diabetes as the outcome and the indicator of food access as the misclassified exposure, a new maximum likelihood estimator (MLE) for Poisson regression with exposure misclassification is needed. The rest of this paper discusses the derivation of the MLE (Section 2), extensive simulations (Section 3), and a case study (Section 4). We conclude by discussing our findings and next steps (Section 5).
@@ 2.1 Model and Notation
We are interested in modeling $Y$, a count outcome offset by $O$, given a binary exposure $X$ and covariate vector $\boldsymbol{Z}$. An error-prone, potentially misclassified version of the exposure, $X^{*}$, is available for all $N$ neighborhoods in the sample, while true $X$ is only available for a subset of $n$ “queried” neighborhoods ($n<N$). We can write the joint probability of a complete (queried) observation as follows:
@@
First, ${\Pr}_{\boldsymbol{\beta}}(Y\mid X,\boldsymbol{Z})$ denotes the Poisson regression for $Y$ given ($X$,$\boldsymbol{Z}$), parameterized by $\boldsymbol{\beta}$. Second, ${\Pr}_{\boldsymbol{\eta}}(X\mid X^{*},\boldsymbol{Z})$ denotes the conditional probability mass function (PMF) of the true $X$ given error-prone $X^{*}$ and additional covariates $\boldsymbol{Z}$. Since $X$ is binary, we estimate this misclassification mechanism using logistic regression, parameterized by $\boldsymbol{\eta}$. To accommodate one-sided misclassification (e.g., only false positives), this model must be modified (Web Appendix A.1) but the rest of the methods remain unchanged. Third, ${\Pr}(X^{*},\boldsymbol{Z})$ denotes the joint marginal probability density or mass function of the error-prone and additional covariates $(X^{*},\boldsymbol{Z})$. Since $(X^{*},\boldsymbol{Z})$ are fully observed, ${\Pr}(X^{*},\boldsymbol{Z})$ drops out of our log-likelihood for $\boldsymbol{\beta}$, so we can leave this model unspecified. Notice that the second equality in (1) assumes conditional independence of ($Y$,$X^{*}$) given ($X$,$\boldsymbol{Z}$) (i.e., surrogacy).
In our setting, all $N$ neighborhoods have an observed $X^{*}$ based on Haversine distances, but only a limited subset of size $n$ have a measurement for $X$ based on true route-based distances. Thus, the data available for a neighborhood depend on whether it was queried:
Observed data =$=\begin{cases}(Y,X,X^{*},\boldsymbol{Z})&\text{ if the neighborhood was queried, and }\\ (Y,X^{*},\boldsymbol{Z})&\text{ otherwise.}\end{cases}$
Queried neighborhoods are assumed to follow (1), while unqueried ones are assumed to follow ${\Pr}_{\boldsymbol{\beta},\boldsymbol{\eta}}(Y,X^{*},\boldsymbol{Z})=\sum\limits_{x=0}^{1}{\Pr}_{\boldsymbol{\beta}}(Y\mid X=x,\boldsymbol{Z}){\Pr}_{\boldsymbol{\eta}}(X=x\mid X^{*},\boldsymbol{Z}){\Pr}(X^{*},\boldsymbol{Z})$ instead.
@@ 2.2 Maximum Likelihood Estimation
×$\times\left\{\sum_{x=0}^{1}{\Pr}_{\boldsymbol{\beta}}(Y_{i}\mid X_{i}=x,\boldsymbol{Z}_{i}){\Pr}_{\boldsymbol{\eta}}(X_{i}=x\mid X_{i}^{*},\boldsymbol{Z}_{i})\right\}^{1-Q_{i}}.$ (2)
The query indicator controls which probability function governs observation $i$ in (2); it evaluates to $1$ for queried neighborhoods (i.e., when $X_{i}$ is available) and to 00 otherwise. Notice that the $N$ neighborhoods are assumed to be independent in this likelihood. Modifying the likelihood to allow spatial autocorrelation is a promising future direction.
The MLEs $\hat{\boldsymbol{\beta}}$ maximize (2) and are found via an EM algorithm (Web Appendix A.2). Standard errors (SEs) for $\hat{\boldsymbol{\beta}}$ are conveniently estimated using numerical derivatives. Our proposed MLEs are implemented in the possum R package, available at www.github.com/sarahlotspeich/possum. Under regularity conditions, missingness at random, and proper specification of both models, the MLEs are asymptotically consistent, normal, and efficient.
@@ 3 Simulations
We conducted simulation studies to test the performance of our MLE against the gold standard, naive, and complete case analyses. The gold standard analysis fit a Poisson regression using $X$ for all observations, while the naive analysis used error-prone version $X^{*}$, and the complete case used $X$ for only the queried subset. Data generation was designed to mimic our data from the Piedmont Triad (Section 4), with primary variations of the sample size $N$, query proportion $q$, and positive predictive value $PPV=\Pr(X=1|X^{*}=1,Z)$.
[…]
We focus here on simulating varied misclassification rates (Section 3.1) and query percentages (Section 3.2). Unless otherwise stated, we fixed $q=0.1$, $PPV=0.6$, $\beta_{1}=0.18$, and outcome prevalence $\approx 10\%$. Additional results with two-sided misclassification in $X^{*}$ and varied outcome prevalence can be found in Web Appendix B.2–B.3. Like in the cases below, the MLE remained unbiased and more efficient than the complete case analysis.
@@ 3.1 Results Under Varied Misclassification Rates
We considered $50\%$ to $90\%$ $PPV$ at sample sizes of $N=390$ and $2200$. The bias of the MLE remained comparable to that of the gold standard and complete case analyses in all settings (Table 1). Even under the most severe misclassification ($PPV=0.5$), the MLE was less than $0.1\%$ biased. However, the naive analysis yielded heavy downward bias ranging from $-8.9\%$ to $-46.5\%$, worsening as $PPV$ decreased and persisting in larger samples.
Table 1: Simulation results for $\hat{\beta}_{1}$, the adjusted log prevalence ratio on $X$, under varied positive predictive value ($PPV$) and sample size ($N$), assuming $\approx 10\%$ outcome prevalence, $\beta_{1}=0.18$, and $q=0.1$ query percentage.
@@
390 0.5 $0.000$ $0.005$ $-0.465$ $0.008$ $0.002$ $0.015$ $0.098$ $0.000$ $0.005$ $0.005$ $0.949$ $0.955$
0.6 $0.000$ $0.005$ $-0.367$ $0.007$ $0.000$ $0.016$ $0.085$ $0.000$ $0.005$ $0.005$ $0.954$ $0.975$
0.7 $0.001$ $0.005$ $-0.271$ $0.007$ $-0.001$ $0.018$ $0.075$ $0.001$ $0.005$ $0.005$ $0.938$ $0.985$
0.8 $0.000$ $0.005$ $-0.181$ $0.007$ $0.000$ $0.019$ $0.059$ $0.000$ $0.005$ $0.005$ $0.953$ $0.985$
0.9 $0.000$ $0.005$ $-0.089$ $0.006$ $0.000$ $0.030$ $0.028$ $0.000$ $0.005$ $0.005$ $0.944$ $0.987$
2200 0.5 $0.000$ $0.002$ $-0.465$ $0.003$ $-0.001$ $0.006$ $0.107$ $0.000$ $0.002$ $0.002$ $0.943$ $0.986$
0.6 $0.001$ $0.002$ $-0.366$ $0.003$ $0.001$ $0.007$ $0.097$ $0.001$ $0.002$ $0.002$ $0.935$ $0.984$
0.7 $0.000$ $0.002$ $-0.272$ $0.003$ $-0.002$ $0.007$ $0.087$ $0.000$ $0.002$ $0.002$ $0.951$ $0.992$
0.8 $0.000$ $0.002$ $-0.179$ $0.003$ $0.000$ $0.008$ $0.065$ $0.000$ $0.002$ $0.002$ $0.948$ $1.002$
0.9 $-0.001$ $0.002$ $-0.089$ $0.003$ $-0.002$ $0.011$ $0.036$ $-0.001$ $0.002$ $0.002$ $0.944$ $0.992$
[…]
The empirical SEs of the MLE were smaller than those of the complete case analysis in every setting. In addition, the average SEs of the MLE approximated the empirical versions very well, which also led to empirical CP in the correct neighborhood of 95%. Finally, the relative efficiency of the MLE to the gold standard analysis was over five times higher than that of the complete case analysis in every setting (between 0.960.960.960.961.001.001.001.00 versus 0.040.040.040.040.110.110.110.11).
@@ 3.2 Results Under Varied Query Percentages
We also considered querying proportions between $q=10\%$ and $75\%$ from samples $N=390$ and $2200$ neighborhoods. The bias of the MLE remained comparable to that of the gold standard analysis; both methods were under 0.5% biased in all settings (Table 2). The empirical SEs of the MLE continued to match or beat those from the complete case analysis in every setting. With the smallest $q$, the SEs from the complete case analysis were over three times the size of the MLE’s. Again, the average SEs of the MLE matched their empirical counterparts, and empirical CP was close to the nominal 95% level.
Table 2: Simulation results for $\beta_{1}$ under varied query percentage ($q$) and sample size ($N$), assuming $\approx 10\%$ outcome prevalence, $\beta_{1}=0.18$, and positive predictive value $=0.6$.
@@
390 0.10 $0.000$ $0.005$ $-0.367$ $0.007$ $0.000$ $0.016$ $0.085$ $0.000$ $0.005$ $0.005$ $0.954$ $0.975$
0.25 $0.002$ $0.005$ $-0.367$ $0.007$ $0.003$ $0.010$ $0.227$ $0.002$ $0.005$ $0.005$ $0.953$ $0.974$
0.50 $0.001$ $0.005$ $-0.366$ $0.007$ $0.002$ $0.007$ $0.463$ $0.001$ $0.005$ $0.005$ $0.953$ $0.992$
0.75 $0.000$ $0.005$ $-0.367$ $0.007$ $0.000$ $0.006$ $0.673$ $0.000$ $0.005$ $0.005$ $0.953$ $1.000$
2200 0.10 $0.001$ $0.002$ $-0.366$ $0.003$ $0.001$ $0.007$ $0.097$ $0.001$ $0.002$ $0.002$ $0.935$ $0.984$
0.25 $0.000$ $0.002$ $-0.368$ $0.003$ $0.000$ $0.004$ $0.227$ $0.000$ $0.002$ $0.002$ $0.949$ $0.980$
0.50 $0.000$ $0.002$ $-0.367$ $0.003$ $0.000$ $0.003$ $0.508$ $0.000$ $0.002$ $0.002$ $0.945$ $0.992$
0.75 $-0.001$ $0.002$ $-0.368$ $0.003$ $-0.001$ $0.002$ $0.695$ $-0.001$ $0.002$ $0.002$ $0.948$ $1.000$
[…]

### v2

@@ 1 Introduction
Maintaining a healthy diet is more than simply a choice. A person must have consistent access to and ability to obtain healthy food. Physical access to healthy foods may be hindered by many factors, including the neighborhood’s geography or community’s socioeconomic makeup (Wolfson et al., 2019). Subsequently, neighborhoods with poor food access may see higher prevalences of diet-related diseases (Gucciardi et al., 2014). For example, Lopez (2007) links food access, specifically the presence of at least one grocery store within a zip code, to a 10% decrease in obesity risk when compared to zip codes without this access. Groups such as food policy councils (Lange et al., 2021) prioritize health equity by supporting policies that increase access to healthy food in rural and minority communities. We focus on the relationship between diabetes prevalence and food access at the neighborhood (census tract) level, although our methods adapt to other outcomes and units as well.
@@
To estimate this relationship, we must first quantify food access, for which there are multiple metrics. We devote our attention to “access indicators,” which classify neighborhoods into binary categories of having/not having access to healthy food within some fixed distance. The United States Department of Agriculture (USDA) includes indicators like these in their Food Research Atlas (U.S. Department of Agriculture, 2019). Computing these food access indicators requires researchers to first calculate the pairwise distances between neighborhoods and potential healthy food retailers. Just like there are multiple ways to measure food access, there are multiple ways to measure these distances. For instance, the Haversine formula modifies the traditional straight-line Euclidean distance between two points (e.g., a neighborhood centroid and grocery store) to account for the curvature of the Earth. Alternatively, route-based distance calculations are possible, which only allow travel via established infrastructure (like roads) to connect the two points.
@@
These methods each have their advantages, and choosing between them to calculate distances requires a trade-off. The Haversine formula is a simple mathematical equation, making it free, easy, and quick to calculate in existing software (e.g., Hijmans, 2022). However, it underestimates distances to healthy food retailers and therefore overestimates neighborhoods’ food access. This underestimation stems from the fact that straight-line distances are necessarily less than or equal to any other measure of distance. In our study, underestimated Haversine distances can lead to neighborhoods being misclassified as having access to healthy foods when they do not (false positives). Specifically, the access indicators are subject to one-sided misclassification; false positives are the only possible type of error.
@@
In short, Haversine distance is easily obtained but leads to access indicators that may be incorrect. Taking route-based distances via software like Google Maps circumvents the possibility of unrealistic, impassable routes. Still, these more-accurate distances are computationally expensive and, despite availability of open source options, may also be costly. Therefore, calculating them for all neighborhoods may not be plausible.
@@
For our study, with diabetes prevalence as the outcome and (potentially mismeasured) food access indicator as the exposure, a new maximum likelihood estimator (MLE) for Poisson regression with exposure misclassification is needed. This MLE incorporates all information for all neighborhoods, offering better statistical efficiency than design-based approaches, like inverse probability weighting. As a model-based approach, it offers flexibility in choosing the queried subset; design-based approaches require nonzero sampling probability for all observations. Moreover, it is computationally faster than multi-stage model-based approaches, like multiple imputation or Bayesian modeling, and it would be straightforward to adapt for more robustness (e.g., semiparametric MLE). The rest of this paper discusses the derivation of the MLE (Section 2), extensive simulations (Section 3), and a case study (Section 4). We conclude by discussing findings and next steps (Section 5).
@@ 2.1 Model and Notation
We model $Y$, a count outcome offset by $O$, given a binary exposure $X$ and covariate vector $\boldsymbol{Z}$. An error-prone, potentially misclassified version of the exposure, $X^{*}$, is available for all $N$ sampled neighborhoods, while true $X$ is only available for a subset of $n$ “queried” ones ($n<N$). We write the joint probability of a complete (queried) observation as follows:
@@
First, ${\Pr}_{\boldsymbol{\beta}}(Y\mid X,\boldsymbol{Z})$ denotes the Poisson regression for $Y$ given ($X$, $\boldsymbol{Z}$), parameterized by log prevalence ratios (log PRs) $\boldsymbol{\beta}$. Second, ${\Pr}_{\boldsymbol{\eta}}(X\mid X^{*},\boldsymbol{Z})$ denotes the conditional probability mass function for $X$ given ($X^{*}$, $\boldsymbol{Z}$). We estimate this misclassification mechanism using logistic regression, governed by $\boldsymbol{\eta}$. To accommodate one-sided misclassification (e.g., only false positives), this model must be modified (Web Appendix A.1) but the rest of the methods remain unchanged. Third, ${\Pr}(X^{*},\boldsymbol{Z})$ denotes the joint probability density or mass function of ($X^{*}$$\boldsymbol{Z}$). Since ($X^{*}$$\boldsymbol{Z}$) are fully observed, ${\Pr}(X^{*},\boldsymbol{Z})$ drops out of our log-likelihood for $\boldsymbol{\beta}$, so we can leave this model unspecified. Notice that the second equality in (1) assumes conditional independence of ($Y$, $X^{*}$) given ($X$, $\boldsymbol{Z}$) (i.e., surrogacy).
In our setting, all $N$ neighborhoods have an observed $X^{*}$ based on Haversine distances, but only a limited subset of size $n$ has true $X$ measured based on route-based distances. Thus, the data available for a neighborhood depend on whether it was queried:
[…]
Queried neighborhoods are assumed to follow (1), while unqueried ones are assumed to instead follow:
${\Pr}_{\boldsymbol{\beta},\boldsymbol{\eta}}(Y,X^{*},\boldsymbol{Z})$ $=\sum_{x=0}^{1}{\Pr}_{\boldsymbol{\beta}}(Y\mid X=x,\boldsymbol{Z}){\Pr}_{\boldsymbol{\eta}}(X=x\mid X^{*},\boldsymbol{Z}){\Pr}(X^{*},\boldsymbol{Z}).$
@@ 2.2 Maximum Likelihood Estimation
[…]
The query indicator controls which probability function governs observation $i$ in (2); it evaluates to $1$ for queried neighborhoods (i.e., when $X_{i}$ is available) and to 0 otherwise. Notice that the $N$ neighborhoods are assumed to be independent in this likelihood. Modifying the likelihood to allow spatial autocorrelation is a promising future direction.
The MLEs $\hat{\boldsymbol{\beta}}$ maximize (2) and are found via an EM algorithm (Web Appendix A.2 and Supplemental Figure S1). Standard errors (SEs) for $\hat{\boldsymbol{\beta}}$ are conveniently estimated using numerical derivatives. Our proposed MLEs are implemented in the possum R package, available at www.github.com/sarahlotspeich/possum. Under regularity conditions, missingness at random, and proper specification of both models, the MLEs are asymptotically consistent, normal, and efficient.
@@ 3 Simulations
We conducted simulation studies to test the performance of our MLEs against the gold standard, naive, and complete case analyses. The gold standard analysis fit a Poisson regression using $X$ for all observations, while the naive analysis used error-prone version $X^{*}$, and the complete case used $X$ for only the queried subset. Data generation was designed to mimic our data from the Piedmont Triad (Section 4), with primary variations of the sample size $N$, query proportion $q$, and positive predictive value $PPV=\Pr(X=1|X^{*}=1,Z)$.
We considered samples of $N=390$ neighborhoods, approximately the number of census tracts in the Piedmont Triad, or $N=2200$ neighborhoods, approximately the number of census tracts in North Carolina. In all settings, we first simulated a covariate ($Z$), followed by a potentially misclassified exposure ($X^{*}\mid Z$), a true exposure ($X\mid X^{*},Z$), an offset ($O$), and a count outcome ($Y\mid X,Z$). Nuisance parameters were chosen to force a prespecified $PPV$. Based on query proportion $q$ ($q\in[0,1]$), a subset of $n=Nq$ neighborhoods had non-missing $X$ for the complete case and MLE analyses. All $n$ were chosen from those with $X^{*}=1$. Since false negatives for food access are not possible, neighborhoods with $X^{*}=0$ must have $X=0$ and thus do not need to be queried. For more details on simulation setup, see Web Appendix B.1. We ran $\mathrm{1000}$ replications per setting and recorded summary metrics capturing empirical bias, variance, and relative efficiency for $\hat{\beta}_{1}$, the adjusted log PR on $X$. We also assessed the empirical coverage probability (CP) of the 95% Wald confidence interval and the validity of the SE estimator.
We focus here on results under varied misclassification rates (Section 3.1) and query percentages (Section 3.2). Unless otherwise stated, we fixed $q=0.1$, $PPV=0.6$, $\beta_{1}=0.18$, and outcome prevalence $\approx 10\%$. Additional results with two-sided misclassification in $X^{*}$ and varied outcome prevalence can be found in Web Appendices B.2–B.3. Like the settings below, the MLE remained unbiased and more efficient than the complete case analysis.
@@ 3.1 Results Under Varied Misclassification Rates
We considered $PPV$ from $0.5$ to $0.9$ at sample sizes of $N=390$ and $2200$. The bias of the MLE remained comparable to that of the gold standard and complete case analyses in all settings (Table 1). Even under the most severe misclassification ($PPV=0.5$), the MLE was less than $0.1\%$ biased. However, the naive analysis yielded heavy downward bias ranging from $9\%$ to $47\%$, worsening as $PPV$ decreased and persisting in larger samples.
Table 1: Simulation results for $\hat{\beta}_{1}$, the adjusted log PR on $X$, under varied positive predictive value (PPVPPV) and sample size (NN), assuming $\approx 10\%$ outcome prevalence, $\beta_{1}=0.18$, and $q=0.1$ query percentage.
@@
390 0.5 $0.000$ $0.005$ 0.465-0.465 $0.008$ $0.002$ $0.015$ $0.098$ $0.000$ $0.005$ $0.005$ $0.949$ $0.955$
0.6 $0.000$ $0.005$ 0.367-0.367 $0.007$ $0.000$ $0.016$ $0.085$ $0.000$ $0.005$ $0.005$ $0.954$ $0.975$
0.7 $0.001$ $0.005$ 0.271-0.271 $0.007$ 0.001-0.001 $0.018$ $0.075$ $0.001$ $0.005$ $0.005$ $0.938$ $0.985$
0.8 $0.000$ $0.005$ 0.181-0.181 $0.007$ $0.000$ $0.019$ $0.059$ $0.000$ $0.005$ $0.005$ $0.953$ $0.985$
0.9 $0.000$ $0.005$ 0.089-0.089 $0.006$ $0.000$ $0.030$ $0.028$ $0.000$ $0.005$ $0.005$ $0.944$ $0.987$
2200 0.5 $0.000$ $0.002$ 0.465-0.465 $0.003$ 0.001-0.001 $0.006$ $0.107$ $0.000$ $0.002$ $0.002$ $0.943$ $0.986$
0.6 $0.001$ $0.002$ 0.366-0.366 $0.003$ $0.001$ $0.007$ $0.097$ $0.001$ $0.002$ $0.002$ $0.935$ $0.984$
0.7 $0.000$ $0.002$ 0.272-0.272 $0.003$ 0.002-0.002 $0.007$ $0.087$ $0.000$ $0.002$ $0.002$ $0.951$ $0.992$
0.8 $0.000$ $0.002$ 0.179-0.179 $0.003$ $0.000$ $0.008$ $0.065$ $0.000$ $0.002$ $0.002$ $0.948$ $1.002$
0.9 0.001-0.001 $0.002$ 0.089-0.089 $0.003$ 0.002-0.002 $0.011$ $0.036$ 0.001-0.001 $0.002$ $0.002$ $0.944$ $0.992$
  • Note: Bias and ESE are, respectively, the empirical relative bias and standard error of each estimator; ASE is the average of the standard error estimator; CP is the empirical coverage probability of the 95% Wald confidence intervals (computed with the standard error estimator); and RE is the relative efficiency of the estimator to the gold standard. All entries are based on $\mathrm{1000}$ replicates. Across all settings, 22 replicates of the MLE ($0.2\%$) were replaced with the naive analysis due to $100\%$ PPV in the queried subset and perfect separation in the misclassification model.
The empirical SEs of the MLE were smaller than those of the complete case analysis in every setting. In addition, the average SEs of the MLE approximated the empirical versions very well, which also led to empirical CP in the correct neighborhood of 95%. Finally, the relative efficiency of the MLE to the gold standard analysis was over five times higher than that of the complete case analysis in every setting (between 0.960.961.001.00 versus 0.040.040.110.11).
@@ 3.2 Results Under Varied Query Percentages
We also considered querying proportions between $q=0.1$ and $0.5$ from samples of $N=390$ and $2200$ neighborhoods. The bias of the MLE remained comparable to that of the gold standard analysis; both methods were under 0.5% biased in all settings (Table 2). The empirical SEs of the MLE continued to match or beat those from the complete case analysis in every setting. With the smallest $q$, the SEs from the complete case analysis were over three times the size of the MLE’s. Again, the average SEs of the MLE matched their empirical counterparts, and empirical CP was close to the nominal 95% level.
Table 2: Simulation results for $\hat{\beta}_{1}$ under varied query percentage ($q$) and sample size (NN), assuming $\approx 10\%$ outcome prevalence, $\beta_{1}=0.18$, and positive predictive value =0.6=0.6.
@@
390 0.10 $0.000$ $0.005$ 0.367-0.367 $0.007$ $0.000$ $0.016$ $0.085$ $0.000$ $0.005$ $0.005$ $0.954$ $0.975$
0.25 $0.002$ $0.005$ 0.367-0.367 $0.007$ $0.003$ $0.010$ $0.227$ $0.002$ $0.005$ $0.005$ $0.953$ $0.974$
0.50 $0.001$ $0.005$ 0.366-0.366 $0.007$ $0.002$ $0.007$ $0.463$ $0.001$ $0.005$ $0.005$ $0.953$ $0.992$
2200 0.10 $0.001$ $0.002$ 0.366-0.366 $0.003$ $0.001$ $0.007$ $0.097$ $0.001$ $0.002$ $0.002$ $0.935$ $0.984$
0.25 $0.000$ $0.002$ 0.368-0.368 $0.003$ $0.000$ $0.004$ $0.227$ $0.000$ $0.002$ $0.002$ $0.949$ $0.980$
0.50 $0.000$ $0.002$ 0.367-0.367 $0.003$ $0.000$ $0.003$ $0.508$ $0.000$ $0.002$ $0.002$ $0.945$ $0.992$
  • Note: Bias and ESE are, respectively, the empirical relative bias and standard error of each estimator; ASE is the average of the standard error estimator; CP is the empirical coverage probability of the 95% Wald confidence intervals (computed with the standard error estimator); and RE is the relative efficiency of the estimator as compared to the gold standard. All entries are based on $\mathrm{1000}$ replicates.